OpenStudy (anonymous):

My answer does not match the calculator on finding the answer of sin(x/2) when tanx=1/2.

hero (hero):

Here's an idea. Post the complete original question. Then post what you believe is the correct answer.

OpenStudy (anonymous):

I'll try to be as consice as possible: The formula for the half angle of sin is: \[\sqrt{\dfrac{1-\cos x}{2}}\]

OpenStudy (anonymous):

I understand the normal procedure of finding the answer but it doesn't make that much sense, in fact, the answer I get from the formula is the answer the calculator gives me for cos(x/2).

OpenStudy (anonymous):

If tanx=1/2 and pi<x<3pi/2, then cosx=-(2sqrt5)/5 and sinx=-(sqrt5)/5.

OpenStudy (anonymous):

Now, by substituting, I would get something like sqrt((1-(-2sqrt5/5))/2) or sqrt((5+2sqrt5)/10)

OpenStudy (anonymous):

Pumas campeón!

OpenStudy (anonymous):

¡Campeón de campeones!